Using logarithmic differentiation to find the derivative of f(x)=[(e^3x√2x-5)/(6-5x)^4]


(2) Find the integral of[ (x^3+1)^1/3x
^5]dx
1
Expert's answer
2020-07-01T15:11:32-0400

Given :


"\frac{(e^{3x})(\sqrt{2x-5)}}{(6-5x)^{4}}=y"


Rewrite as "{(lne^{3x})+ln\sqrt{(2x-5)}}-{ln(6-5x)^{4}}=lny"


"{(3x})+{\frac{1}{2}*{ln(2x-5)}}-{4ln (6-5x)}=ln y"

Differentiate


3+"\frac{1}{2x-5}-\frac{4(-5)}{6-5x} = \frac{1 y'}{y}"

"y'= y(3+\frac{1}{2x-5}+\frac{20}{6-5x})"


Solution : y' = e3x"√2x-5)/(6-5x)^4](3+\frac{1}{2x-5}+\frac{20}{6-5x})"



Part 2:


To find : "\int\frac{\left(1+x^{\left(3\right)^{ }}\right)^{\frac{1}{3}}}{x^{5}} dx"


solution : Take u= "\frac{\left(1+x^{\left(3\right)^{ }}\right)^{\frac{1}{3}}}{x^{ }}"


du= "\frac{x}{\left(1+x^{\left(3\right)}\right)^{\frac{2}{3}}}-\frac{\left(1+x^{\left(3\right)^{ }}\right)^{\frac{1}{3}}}{x^{2}}" dx


Hence "\int\frac{\left(1+x^{\left(3\right)^{ }}\right)^{\frac{1}{3}}}{x^{5}} dx" = "\int" -u3 du

= -"\frac{u^{4}}{4} +c"


substitute for u , and the solution is =- "\frac{(x^{3}+1)^{\frac{4}{3}}}{4x^{4}} +c"